AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 15 of 33  ELOG logo
ID Date Author Subject
  381   Tue Nov 29 14:37:56 2022 AR TDTuesday 29 November

16:20

DSSD bias and leakage current values:

-99.99 V

-13.2535 uA

Pulsar peak widths (ch FWHM)

aida01 ~ 100

aida02 ~ 103

aida03 ~ 128

aida04 ~ 133

aida05 ~ 22

Pretty similar to what was found in august https://elog.ph.ed.ac.uk/CARME/377

 

 

 

 

Attachment 1: IMG_8240_3.jpg
IMG_8240_3.jpg
Attachment 2: IMG_8241_3.jpg
IMG_8241_3.jpg
Attachment 3: IMG_8242_3.jpg
IMG_8242_3.jpg
Attachment 4: IMG_8243_3.jpg
IMG_8243_3.jpg
Attachment 5: IMG_8246_3.jpg
IMG_8246_3.jpg
Attachment 6: IMG_8247_3.jpg
IMG_8247_3.jpg
Attachment 7: IMG_8248_3.jpg
IMG_8248_3.jpg
Attachment 8: IMG_8249_3.jpg
IMG_8249_3.jpg
Attachment 9: IMG_8250_3.jpg
IMG_8250_3.jpg
Attachment 10: IMG_8252_3.jpg
IMG_8252_3.jpg
Attachment 11: IMG_8253_3.jpg
IMG_8253_3.jpg
Attachment 12: IMG_8254_3.jpg
IMG_8254_3.jpg
Attachment 13: IMG_8255_3.jpg
IMG_8255_3.jpg
Attachment 14: IMG_8257_3.jpg
IMG_8257_3.jpg
Attachment 15: IMG_8258_3.jpg
IMG_8258_3.jpg
Attachment 16: IMG_8259_3.jpg
IMG_8259_3.jpg
Attachment 17: IMG_8264_3.jpg
IMG_8264_3.jpg
Attachment 18: Screenshot_2022-11-29_at_16.40.13.png
Screenshot_2022-11-29_at_16.40.13.png
Attachment 19: Screenshot_2022-11-29_at_16.42.37.png
Screenshot_2022-11-29_at_16.42.37.png
Attachment 20: Screenshot_2022-11-29_at_16.47.03.png
Screenshot_2022-11-29_at_16.47.03.png
Attachment 21: Screenshot_2022-11-29_at_16.49.07.png
Screenshot_2022-11-29_at_16.49.07.png
Attachment 22: Screenshot_2022-11-29_at_16.49.44.png
Screenshot_2022-11-29_at_16.49.44.png
  380   Mon Nov 28 13:14:33 2022 AR TDMonday 28 November
Attachment 1: 91F62DFF-B008-4204-91E5-BD041D8E4254.jpeg
91F62DFF-B008-4204-91E5-BD041D8E4254.jpeg
Attachment 2: EA4AFB0D-7960-4C5D-834D-AA8B5932DAC2.jpeg
EA4AFB0D-7960-4C5D-834D-AA8B5932DAC2.jpeg
Attachment 3: F05ACF42-F97E-4F0F-9174-0B997A86136F.jpeg
F05ACF42-F97E-4F0F-9174-0B997A86136F.jpeg
Attachment 4: 02F9137F-A28C-4369-B0DB-EEBCEC2A5F5C.jpeg
02F9137F-A28C-4369-B0DB-EEBCEC2A5F5C.jpeg
Attachment 5: 9ED571DA-3088-43D6-B186-44ABEA96DB93.jpeg
9ED571DA-3088-43D6-B186-44ABEA96DB93.jpeg
Attachment 6: F3526AC5-E4B8-4D8A-9D88-57B4A46B4B0D.jpeg
F3526AC5-E4B8-4D8A-9D88-57B4A46B4B0D.jpeg
Attachment 7: 80088713-1F6E-4AAB-AEA0-77E6FC78DAF9.jpeg
80088713-1F6E-4AAB-AEA0-77E6FC78DAF9.jpeg
Attachment 8: 9B5BCDB9-6774-4873-A5BD-FC2960FB3FFF.jpeg
9B5BCDB9-6774-4873-A5BD-FC2960FB3FFF.jpeg
Attachment 9: 9340303A-4F43-4C31-B420-3A4D6ACF4EA7.jpeg
9340303A-4F43-4C31-B420-3A4D6ACF4EA7.jpeg
Attachment 10: 574D2470-3B5E-499E-A9F7-C561E8535115.jpeg
574D2470-3B5E-499E-A9F7-C561E8535115.jpeg
Attachment 11: 227ED3B1-7F1E-4529-90C3-CE84CAFA121C.jpeg
227ED3B1-7F1E-4529-90C3-CE84CAFA121C.jpeg
Attachment 12: CB94FBF3-2FFF-4795-B03D-5FD8CFE34242.jpeg
CB94FBF3-2FFF-4795-B03D-5FD8CFE34242.jpeg
  379   Mon Oct 3 16:03:33 2022 OH, CBTTT14 Photos
Photos of the new TTT14 DSSD delivered to GSI.
Photos of detectors are broken up by photos of SN on box
To big to put on elog. Have uploaded to /Disk/ds-sopa-group/np/GSI/CARME/TTT14_Photos.zip and will change permissions to 777
  378   Sun Aug 14 06:50:58 2022 TD, JMSunday 14 August
07.51 DAQ continues files 22Feb/R97

      ASIC settings 2022Jan31-11-52-00
       slow comparator 0x64

      EG&G Ortec 448 Pulser OFF

      All system wide checks OK

Attachment 1 - ADC data item stats

Attachment 2 - FEE64 temps OK

Attachments 3 & 4, 8 & 9 - per FEE64 1.8.W spectra 20us FSR and 200us FSR

Attachments 5 & 6 - per FEE64 1.8.L spectra

Attachment 7 - per FEE64 stat spectra

Leakage current remains 8.36 uA, 100V. 

Bias power OFF, SIP ion pump power back ON. NIM crate powered off. Motors pi and nnrpi1 still on for remote access. 
Compressed air off, water off. Fee relay power off and power cable disconnected. Motors box off. Safety pin for motors reinserted. 

Checked cable harnesses at cryring. 2 good short sections found and left with boxes in cryring. 2 damaged long sections (bad pin and broken ceramic) taken to repair in 
Edinburgh.
Attachment 1: 60.PNG
60.PNG
Attachment 2: 59.PNG
59.PNG
Attachment 3: 58.PNG
58.PNG
Attachment 4: 57.PNG
57.PNG
Attachment 5: 56.PNG
56.PNG
Attachment 6: 55.PNG
55.PNG
Attachment 7: 54.PNG
54.PNG
Attachment 8: 61.PNG
61.PNG
Attachment 9: 62.png
62.png
Attachment 10: 20220814_091209.jpg
20220814_091209.jpg
  377   Sat Aug 13 10:23:59 2022 JM, TDNoise tests pt3 and pulser walk

Waveforms and discriminator ON. Pulser re-connected to all fees via daisy chain. V=100v, I=8.36uA

 

Pulser peak widths (ch FWHM)

aida01 ~ 109

aida02 ~ 98

aida05 ~ 24

aida03 ~ 118

aida04 ~ 136

Attachments 2, 3 & 4 1.8.W spectra, 20us FSR

Attachment 5 FEE64 temps OK

Attachment 6 per FEE64 ADC data item rates OK

Attachment 7 per FEE64 rate spectra

 

Merger reset following procedure from https://elog.ph.ed.ac.uk/DESPEC/36. Keywords for future elog search NetVar, xfer links

 

 

Pulser walkthrough R94

EG&G Ortec 448 Pulser settings

Amplitudes 90,000 - 10,000 @ 10,000 step

Attenuators x5 IN

t_r 50ns tau_d 50us

polarity + (- polarity via Cooknell SA1)

Frequency int relay

Motors box turned ON, motors pi ON, dlink network connector ON. Motors box, pi and dlink all in 19 inch rack but power supply is from a uk mains extender daisy chained to vacuum mains extender.  19 inch rack power supply remains on clean line.

Pulser peak widths (ch FWHM)

aida01 ~ 105

aida02 ~ 98

aida05 ~ 23

aida03 ~ 132

aida04 ~ 144

See attachments 8 & 9

 

Detectors moved in by pnuematic only. Leakage current increased to 9.6 uA (same as leakage current during bias ramping)

Pulser Walk R95

Pulser settings as above

See attachments 10 & 11

 

Detectors moved back out . Motors box, pi and dlink still turned ON. Leakage current to 8.36 uA

Pulser Walk R96

Pulser settings as above

See attachments 1 & 12

 

 

Jumper removed from nn bias fee adaptor board (aida04). No jumpers on any adaptor boards now. V=100V, I=8.36uA

See attachments 14 & 15 - 1.8.W spectra 20us FSR

Pulser on 90,000 x5 attenuate

aida01 ~ 123, centroid ~14691

aida02 ~ 106, centroid ~ 15038

aida05 ~ 23, centroid ~ 14414

aida03 ~ 125, centroid ~ 50838

aida04 ~ 127, centroid ~ 51754

 

Jumper reinstalled on aida04 nn bias. Note only link 1, link2 inaccesible with fee inserted into ERNI. See attachment 13

aida01 ~ 123, centroid ~14688

aida02 ~ 114, centroid ~ 15034

aida05 ~ 23, centroid ~ 14411

aida03 ~ 118, centroid ~ 50854

aida04 ~ 135, centroid ~ 51697

 

Pulser off. Background run to be left overnight started 14:24 R97. V=100v, I=8.36 uA.

Attachment 1: 51.PNG
51.PNG
Attachment 2: 40.PNG
40.PNG
Attachment 3: 41.PNG
41.PNG
Attachment 4: 42.PNG
42.PNG
Attachment 5: 43.PNG
43.PNG
Attachment 6: 44.PNG
44.PNG
Attachment 7: 45.PNG
45.PNG
Attachment 8: 46.PNG
46.PNG
Attachment 9: 47.PNG
47.PNG
Attachment 10: 48.PNG
48.PNG
Attachment 11: 49.PNG
49.PNG
Attachment 12: 50.PNG
50.PNG
Attachment 13: 20220811_172335.jpg
20220811_172335.jpg
Attachment 14: 53.PNG
53.PNG
Attachment 15: 52.PNG
52.PNG
  376   Fri Aug 12 15:12:37 2022 JM, TDCARME measurements

measurement 1 -string
intermediate (smaller) chamber - 2144 (mm) circumference,  measured half circumference 1072 mm 
outer (larger) chamber - 2160 (mm) circumference, measured 1/4 circumference 540mm

measurement 2 - foil tape
intermediate (smaller) chamber - 2156 (mm) circumference,  measured half circumference 1078 mm 
outer (larger) chamber - 2172 (mm) circumference, measured 1/4 circumference 543mm

Difference from centre to flat surface on outer chamber is 344 mm on drawing -> say 346 mm radius -> 2174 mm circumference
Part 620FBM600 from pfeiffer intermediate chamber 690mm diameter, 2168mm 

Measurements of parts of CARME frame to the target frame as a reference for target frame position.

Attachments 3+4 show clearance between CARME frame and target frame just below interaction chamber from inside and outside of the ring. Attachment 5 is clearance in same section but above interaction chamber

Attachment 6 - measurement between outer edge of CARME frame and inner edge of target frame - 165mm

Attachment 7 - measurements between outer edge of CARME frame and inner edge of target frame - 455 mm 

Attachment 8 - measurement between lower edge of CARME frame and lower edge of target frame - 1245 mm

Attachment 9 - measurement between top of VAT valve and lower edge of target frame - 233mm (left arrow), 216mm (right arrow)

 

Attachment 1: 20220810_095036.jpg
20220810_095036.jpg
Attachment 2: 20220810_111638.jpg
20220810_111638.jpg
Attachment 3: 20220813_142817.jpg
20220813_142817.jpg
Attachment 4: 20220813_142930.jpg
20220813_142930.jpg
Attachment 5: 20220813_142945.jpg
20220813_142945.jpg
Attachment 6: measurement1.png
measurement1.png
Attachment 7: measurement2.png
measurement2.png
Attachment 8: measurement3.png
measurement3.png
Attachment 9: measurement4.png
measurement4.png
  375   Fri Aug 12 09:06:57 2022 TD, JMNoise tests pt2
10.07 DSSSD bias -100V leakage current -8.65uA

Waveforms and discriminators off, still on clean power supply

stats are ~10k across all fees

aida03 ~ 220
aida04 ~ 205
aida01 ~ 378
aida02 ~ 259
aida05 ~ 40

Grounding removed from all fees except nn bias fee aida04 and aida05. Grounding goes into both adaptor board and fee body

aida01 - no identifiable peak
aida02 - no identifiable peak
aida05 - double peak separated by 70 chn
aida04 - 400 shouldering
aida03 - no identifiable peak

Grounding removed from all fees except nn and pn bias fees aida04, aida01 and aida05. Grounding goes into both adaptor board and fee body

aida01 - 405
aida02 - no identifiable peak
aida05 - double peak
aida04 - 228
aida03 - 417

Grounding removed from all fees except pn bias fee aida01 and aida05. Grounding goes into both adaptor board and fee body

aida01 - 383
aida02 - no identifiable peak
aida05 - 64
aida04 - no identifiable peak
aida03 - no identifiable peak

Grounding on all fees. Grounding goes into adaptor board only

aida01 - no identifiable peak
aida02 - no identifiable peak
aida05 - 63  (wasn't changed)
aida04 - 700
aida03 - 700

Back to previous grounding on all fees

aida01 - 331
aida02 - 226
aida05 - 37
aida04 - 207
aida03 - 214

power cycle

aida01 - 326
aida02 - 238
aida05 - 46
aida04 - 212
aida03 - 195

Waveforms and discriminator back on

Removed test inputs and terminators from all fees, except aida03 (unable to reach) however the lemo is not touching the chamber
See attached waveforms compared to previous elog 32,33

Inserted test input to aida05 only directly from pulser. Width ~ 28

Inserted test input to aida01 only directly from pulser. Width ~ 329

Waveforms and discriminator off -> no impact on aida01 pulser width

Waveforms and discriminator back on

Connected a grounding cable from one of the water cooling pipes to cryring ground

aida01 ~ 205
aida05 ~ 23
aida02 ~ 105
aida03 ~ 71
aida04 ~ 148
Fees with bias attached to them aida01 (+ve) and aida04 (-ve) both show larger widths than there polarity counterpart
waveform screencap 34,35

Fee probe
7.46V left most bar
-6.26V 2nd bar
5.26V 3rd bar
5.26V 4th bar
5.25V 5th bar

Power supply probe
Top -6.05
2nd -6.05
3rd -6.05
4th +6.49
5th -7.78

new Ortec 448 pulser swapped in
aida04 ~ 166
aida05 ~ 21

Attachment 7-11 oscilloscope picture when probing fee relay power supply. Power ON fees connected. Top is attachment 7, bottom (5th) is attachment 11.  
Attachment 12-16 old pulser 
Attachment 17-18 new pulser
Attachment 19-23 oscilloscope picture when probing fee relay power supply. Power OFF fees connected. Top is attachment 19, bottom (5th) is attachment 23.
Attachment 24-28 oscilloscope picture when probing fee relay power supply. Power ON, fees disconnected. Top is attachment 24, bottom (5th) is attachment 28.
Attachment 29,30 probe-scope setup    
Attachment 1: 33.PNG
33.PNG
Attachment 2: 32.PNG
32.PNG
Attachment 3: 31.PNG
31.PNG
Attachment 4: 30.PNG
30.PNG
Attachment 5: 34.PNG
34.PNG
Attachment 6: 35.PNG
35.PNG
Attachment 7: 20220812_163204.jpg
20220812_163204.jpg
Attachment 8: 20220812_163230.jpg
20220812_163230.jpg
Attachment 9: 20220812_163319.jpg
20220812_163319.jpg
Attachment 10: 20220812_163339.jpg
20220812_163339.jpg
Attachment 11: 20220812_163411.jpg
20220812_163411.jpg
Attachment 12: 20220812_163526.jpg
20220812_163526.jpg
Attachment 13: 20220812_163601.jpg
20220812_163601.jpg
Attachment 14: 20220812_163650.jpg
20220812_163650.jpg
Attachment 15: 20220812_163705.jpg
20220812_163705.jpg
Attachment 16: 20220812_163744.jpg
20220812_163744.jpg
Attachment 17: 20220812_165322.jpg
20220812_165322.jpg
Attachment 18: 20220812_165330.jpg
20220812_165330.jpg
Attachment 19: 20220812_165619.jpg
20220812_165619.jpg
Attachment 20: 20220812_165639.jpg
20220812_165639.jpg
Attachment 21: 20220812_165652.jpg
20220812_165652.jpg
Attachment 22: 20220812_165704.jpg
20220812_165704.jpg
Attachment 23: 20220812_165717.jpg
20220812_165717.jpg
Attachment 24: 20220812_165812.jpg
20220812_165812.jpg
Attachment 25: 20220812_165839.jpg
20220812_165839.jpg
Attachment 26: 20220812_165849.jpg
20220812_165849.jpg
Attachment 27: 20220812_165900.jpg
20220812_165900.jpg
Attachment 28: 20220812_165911.jpg
20220812_165911.jpg
Attachment 29: 20220812_170337.jpg
20220812_170337.jpg
Attachment 30: 20220813_113236.jpg
20220813_113236.jpg
  374   Thu Aug 11 09:25:13 2022 JM, TDNoise tests


Moved pneumatic motor in and out several revolutions

Biased detector up to 100 V, I ~ 12.4 uA ambient temp c. 21 deg C

Moving detector in increased leakge current to ~14.5 uA, moving detector out dropped leakage current back to ~12.4 uA

aida05 - pulser width 21 chn (slightly lower than yesterday)

aida01 - pulser width 124 chn

aida02 - double peak (is not currently grounded)

aida03 - 183 chn

aida04 - 212 chn

 

Attachment 1     1.8.L spectra p+n FEE64s

Attachment 2      per FEE64 rate spectra

Attachment 3      aida05 1.8.L, 1.8.W and rate spectra

Attachment 4      FEE64 temperatures - OK

Attachment 5      per FEE64 ADC data event stats - rates for aida01-aida04 significabtly lower than yesterday presumably due to detector bias -100V

Attachments 6 & 7 per FEE64 1.8.W spectra

Attachment 8      1.8.L spectra for n+n FEE64s

regrounded aida02 back to cryring ground - required removing star configuration of all fee grounds and making a new

aida05 - pulser width 30 chn 

aida01 - pulser width 300 chn

aida02 - 220 -> double peak gone 

aida03- 190

aida04 - 200

Perhaps star configuration isn't as tight as before - re tighten star section

Re tightened section -> no impact on pulser widths

Moved detector in using pneumatic -> no impact on pulser widths

Added a grounding cable from the 19 inch rack to the same CRYRING ground that the fees are connected too. Stats around 2-3k, slow comparator = 0x64, bias=100v, I=12.18 uA

aida01  ~ 260

aida02 ~ 177

aida03 ~ 178

aida04 ~ 187

aida05 ~28

Removed CRYRING ground. Fees and rack connected together still. 

no change

Bias supply module is floating

ion pumps off, gauge off -> no change in waveforms, leakage current dropped to 8.7 uA

aida01 ~ 255

aida02 ~ 176

aida05 ~ 26

Removed all vacuum systems and motors power from 19 inch rack power. Connected to the rack are now: NIM bin, network switch, 2 inputs to mains relay, raspberrypi, pi dlink usb connector, VME crack

aida01 ~ 220

aida02 ~ 154

aida05 ~ 27

aida03 ~ 181

aida04 ~ 193

Moved power cable for 19 inch rack to CRYRING 'clean' power source

aida03 ~ 211

aida04 ~ 212

aida01 ~ 317

aida02 ~ 215

aida05 ~ 29

Reconnected CRYRING ground 

aida01 ~ 301

aida02 ~ 203

aida05 ~ 28

aida03 ~ 181

aida04 ~ 191

Current status 17:00:
All CARME vacuum equipment is off (turbo, SIP power modules, NEG power modules, gauges) - roughing pump is on but is valved off from the turbo which is valved off from chamber

Motors and motors pi are off. Detectors are fully out and the pin has been reinserted. 

Bias ON 100V, 8.5 uA. Pulser ON, all equipment connected to 19 inch rack power ON.  Jumper is on link 1 for nn bias adaptor card only. 

Attachment 1: 19.PNG
19.PNG
Attachment 2: 18.PNG
18.PNG
Attachment 3: 17.PNG
17.PNG
Attachment 4: 16.PNG
16.PNG
Attachment 5: 15.PNG
15.PNG
Attachment 6: 14.PNG
14.PNG
Attachment 7: 13.PNG
13.PNG
Attachment 8: 12.PNG
12.PNG
Attachment 9: 11.PNG
11.PNG
Attachment 10: 10.PNG
10.PNG
Attachment 11: 9.PNG
9.PNG
Attachment 12: 8.PNG
8.PNG
Attachment 13: 7.PNG
7.PNG
Attachment 14: 6.PNG
6.PNG
  373   Wed Aug 10 10:33:03 2022 TD, ML, JM, NPMeeting at YR09 to discuss options for CARME operations to February 2022
Initial discussions of proposals to provide open CARME chamber and provide access to near and far detector chambers

We assume that we need to complete all work installing and testing DSSSDs and re-assembling CARME by Feb 2022 in readiness for bakeout. 




Proposal #1

Disconnect far detector chamber section from mid-chamber section
Disconnect DN350CF flange from near detector chamber 

Crane to support near detector chamber and mid-chamber sections

Remove feet of near detector chamber and mid-chamber sections

Install cradle(s) beneath near detector chamber and mid-chamber sections

Use for fork lift to remove cradle(s) + near detector chamber and mid-chamber sections - ML confirms that an appropriate fork lift is available 


Proposal #2 (suggested by ML)

Design and fabricate table plus rails to extend transverse rails of CARME chamber support assembly
The CARME chamber support assembly comprises the 'feet' plus two-stage welded Al frame 

Disconnect far detector chamber section from downstream section
Disconnect DN350CF flange from near detector chamber 

Slide CARME chamber plus upper stage of Al support frame transversely onto table

Use crane to move CARME chamber to work area


Other discussions

Discussed workarounds to enable install of FEE64s (outside ring, upper feedthrough) which currently conflicts with longitudinal beam of gas jet target support frame.


To Do

Edinburgh

- investigate feasibility of proposal #2
- investigate how much space is required to install FEE64s (outside ring, upper feedthrough)
  e.g. by extension of transverse beams of gas jet target support frame, move current assembly outwards from ring c. 20-25cm


GSI

- investigate feasibility of gas jet target support frame changes outlined above e.g. movement of gas/water services, purchase of longer cross beams etc.


Suggest meeting (online) to discuss results of these investigations and continue discussions w/c August 22
Attachment 1: 20220810_111638.jpg
20220810_111638.jpg
Attachment 2: 20220810_095036.jpg
20220810_095036.jpg
Attachment 3: 20220810_095150.jpg
20220810_095150.jpg
Attachment 4: 20220810_095159.jpg
20220810_095159.jpg
  372   Wed Aug 10 10:30:04 2022 TD, JMWednesday 10 August
09.45 Found CARME electrical power off

      Reset power switch of distribution box to on
      Reset fuse of 19" rack extender board

      One of the SIP power modules may be faulty - others switch on OK


14.00 Checked coolant water flow OK

      Transferred Netgear 24 port Gbit switch from AIDA (S4) to CARME (YR09)

      Re-connected FEE64 RJ45 network cables
      Re-connected nnrpi1 RJ45 network cable
      Re-connected RJ45 network cable (yellow) to bottom left connector of YR09 patch panel 
      
      @CRYRING user control rooms

      Connected RJ45 netwwork cable (yellow) from bottom left connector of patch panel (labelled "1-2->YR10" and "3-4->YR10") to CARME server carme-gsi p4p1 port


Retarted HttpdTcl servers on nnrpi1 and carme-gsi.

Normal 5x FEE64 power up

All system wide checks OK

FEE64 temperatures OK - attachment 1

ASIC settings 
 slow comparator 0x64

Pulser running through aida01-05
Detectors not biased -> lot of noise aida01-04

AIDA stats - attachments 3 & 4

aida05 pulser width ~150 channels FWHM, normal settings, shaping time 8us - attachment 4

See attachments 2-4

Tried changing ADC control register to 0xf in local controls tab -> stops waveform producing data -> no effect on pulser width (slight increase to 160 chn)
fast comparator blanked off (enable disable mask in discriminator control) -> no effect on pulser width
aida05 was ungrounded -> grounded aida05 to the same ground as aida01-04 -> pulser width reduced to 46 chns
Set ADC control register back to 0x0 -> waveforms back -> pulser width reduced to 24 chns
Changed shaping time from 8ms to 1ms -> pulser width increased to 67 chns -> changed back to 8ms shaping time -> returns to 24 chn width
Changed shaping time from 8ms to 2ms -> pulser width increased to 52 chns
Changed shaping time from 2ms to 4ms -> pulser width decreased to 33 chns
Attachment 1: 4.PNG
4.PNG
Attachment 2: 3.PNG
3.PNG
Attachment 3: 2.png
2.png
Attachment 4: 1.png
1.png
  371   Tue Aug 9 16:07:33 2022 JM, TDAnydesk setup command line

To log into appc218 desktop:
same login and password as to carme@atppc022.gsi.de

To start anydesk

1. make sure a monitor is plugged into the carme computer!
2. ssh into npg@carme-gsi
3. su terminal
4. killall anydesk
5. systemctl start anydesk
returns an anydesk backend, service and tray
Anydesk is now accesible 
 

  370   Tue Aug 9 13:44:16 2022 TD, JMTuesday 9 August
14.00 CARME vacuum c. 1.8e-10 mbar - see attachment 1

open water inlet/outlet taps

NIM bin (MACBs, Ortec 480 Pulser, CAEN N1419ET) switched ON

USB-controlled ac mains relay switched ON - water & interlock LEDs active

Require VETAR2 (previous returned to Nik Kurz) and network switch (currently with AIDA@DESPEC)

Turbo pump ON - normal operation

Ion pumps ON - normal operation

16.30 ML informs us that ambient temperature of CRYRING cave has been increased from c. 16 deg C to c. 21 deg C and that this is likely the cause of the change in vacuum 
pressure in YR09 and the rest of the storage ring.

obtained VETAR2 and crate 

anydesk restarted
Attachment 1: Capture_1.PNG
Capture_1.PNG
Attachment 2: 20220809_163450.jpg
20220809_163450.jpg
Attachment 3: 20220809_135840.jpg
20220809_135840.jpg
Attachment 4: 20220809_135833.jpg
20220809_135833.jpg
Attachment 5: 20220809_135801.jpg
20220809_135801.jpg
  369   Thu Mar 3 15:14:21 2022 JM, OHCARME packing up

Pulser walkthrough taken for motor position 23.1mm and motor position 67.43mm.

Points taken at 30,50,70,90,110,130,150,170,190,210,230 for 23.1mm position and 230,210,190,170,150,130,110 for 67.43mm position.

Motors moved to fully OUT position and pin put back in. Current positions before closing program 19.1mm and 19.57mm. 

Fee power OFF. Fee power module switched OFF and AC mains cable unplugged.  

Bias suppply OFF. NIM crate for bias module switched OFF. 

Water supply OFF

CARME ion gauge turned back ON

  368   Wed Mar 2 22:29:48 2022 CB, SDLast shift with beam

22:30 Shift handed over. DAQ OK. Target OK. Beam OK.

 

00:22 Temperatures OK

Stats OK

Good events OK

Scalers OK

System wide checks as usual - WR fails by 68 ns

Leakage current OK

Target OK

Motors OK

Pressure in ring 5E-10 mbar OK

Beam OK

 

02:26 - All OK as above

WR now fails by 80

Data rate ~800 KB/s

 

03:40 All OK as above

Data rate back to ~ 1900 KB/s

 

04:20 All OK as above

Started compressing files in WS3, monitor 1

Noticed bias is not as stable as it has been so far, but no reason for concern. See attach.

 

04:30 Saw a spike in counts similar to beam halo. See attachment. Did not see it before. No effects on leakage current. To be checked offline.

 

06:00 Could not see any more spikes

All other checks OK as before.

Data rate back to ~1000 Kb/s

 

06:32 Requested beam off for end of beamtime.

Left DAQ on - run R88

Turning off jet target as per Nikos instructions

Final cleanup + pulse walkthrough to be done by JM and OH later today.

 

Target nozzle removed - pressure in S1 now 1.93E-8 mbar

Valves closed

Entering cave to switch compressor off and change N backing pressure to 10 bar.

Attachment 1: Stats_2022-03-03_00-23-00.png
Stats_2022-03-03_00-23-00.png
Attachment 2: Scalers_2022-03-03_00-24-03.png
Scalers_2022-03-03_00-24-03.png
Attachment 3: GoodEvents_2022-03-03_00-23-38.png
GoodEvents_2022-03-03_00-23-38.png
Attachment 4: Ruth_2022-03-03_00-25-48.png
Ruth_2022-03-03_00-25-48.png
Attachment 5: Leakage_022-03-03_00-26-44.png
Leakage_022-03-03_00-26-44.png
Attachment 6: Leakage_2022-03-03_04-22-52.png
Leakage_2022-03-03_04-22-52.png
Attachment 7: ExTotal_2022-03-03_04-27-38.png
ExTotal_2022-03-03_04-27-38.png
Attachment 8: EyTotal_2022-03-03_04-27-10.png
EyTotal_2022-03-03_04-27-10.png
Attachment 9: Spike_2022-03-03_04-29-57.png
Spike_2022-03-03_04-29-57.png
  367   Wed Mar 2 21:16:06 2022 JM, PMCurrent status

22:16

Temperature setpoint = 113k

Not observed noise burst again

Target density OK, S1 between 2E-8 and 3E-8mbar, E4 between 1E-9 and 1E-8 mbar.

Detector moving well, potentionmer between 23.1 mm and 67.43 mm, move in/move out counters the same. Correlation scalar on aida1

Ruthrford peak varies between 10^3 and 10^2 peak height depending on detector position, noise rate also varies with position.

Beam intensity ~1E8

Merger OK

Data rate ~ 1200 kb/sec

Temperatures OK

System checks OK - wr fails

         Base         Current     Difference
aida01 fault      0xab60 :      0xaba4 :      68  
aida02 fault      0x88d3 :      0x8915 :      66  
aida03 fault      0x23aa :      0x23ee :      68  
aida04 fault      0x15d7 :      0x161b :      68  
aida05 fault      0x2f03 :      0x2f20 :      29  
White Rabbit error counter test result: Passed 0, Failed 5

Understand the status reports as follows:-
Status bit 3 : White Rabbit decoder detected an error in the received data
Status bit 2 : Firmware registered WR error, no reload of Timestamp
Status bit 0 : White Rabbit decoder reports uncertain of Timestamp information from WR

Ring pressure oscillation is 30 minute cycle on all areas of the ring.

  366   Wed Mar 2 19:25:25 2022 JM, PMCurrent status

20:25

Observed short burst of low energy noise on aida1, lasted <30s (attachment 8).

Temperature setpoint = 113 K

Target density OK S1 vaies between 2E-8 and 3E-8 mbar, E4 between 1E-9 and 1E-8 mbar.

Detector moving well between 23.1 mm and 67.43mm, move in and move out counters the same. Correlation scalar on aida1.

Rutherford peak varies between 10^3 and 10^2 depending on detector position, noise peak varies with detector position -> lower further in.

Beam intensity ~1E8

Merger OK

Data rate ~ 1900 Kb/sec

Leakage current = 3.208 uA

Temperatures OK

System checks - wr fails

         Base         Current     Difference
aida01 fault      0xab60 :      0xaba4 :      68  
aida02 fault      0x88d3 :      0x8915 :      66  
aida03 fault      0x23aa :      0x23ee :      68  
aida04 fault      0x15d7 :      0x161b :      68  
aida05 fault      0x2f03 :      0x2f20 :      29  
White Rabbit error counter test result: Passed 0, Failed 5

Understand the status reports as follows:-
Status bit 3 : White Rabbit decoder detected an error in the received data
Status bit 2 : Firmware registered WR error, no reload of Timestamp
Status bit 0 : White Rabbit decoder reports uncertain of Timestamp information from WR

Avg ring pressure = 5.5 E-10 mbar. Ring pressure tends to oosciilate between ~4 and 5.5 E-10 mbar

21:21

Burst of noise not observed again

Temperature setpoint = 113 K

Target density OK S1 vaies between 2E-8 and 3E-8 mbar, E4 between 1E-9 and 1E-8 mbar.

Detector moving well between 23.1 mm and 67.43mm, move in and move out counters the same. Correlation scalar on aida1.

Rutherford peak varies between 10^3 and 10^2 depending on detector position, noise peak varies with detector position -> lower further in.

Beam intensity ~1E8

Merger OK

Data rate ~ 1900 Kb/sec

Leakage current = 3.206 uA

Temperatures OK

System checks - wr fails

         Base         Current     Difference
aida01 fault      0xab60 :      0xaba4 :      68  
aida02 fault      0x88d3 :      0x8915 :      66  
aida03 fault      0x23aa :      0x23ee :      68  
aida04 fault      0x15d7 :      0x161b :      68  
aida05 fault      0x2f03 :      0x2f20 :      29  
White Rabbit error counter test result: Passed 0, Failed 5

Understand the status reports as follows:-
Status bit 3 : White Rabbit decoder detected an error in the received data
Status bit 2 : Firmware registered WR error, no reload of Timestamp
Status bit 0 : White Rabbit decoder reports uncertain of Timestamp information from WR

Avg ring pressure 5.4E-10 mbar

Attachment 1: 2022-03-02_20-37-20-rateIN.png
2022-03-02_20-37-20-rateIN.png
Attachment 2: 2022-03-02_20-36-42-rateOUT.png
2022-03-02_20-36-42-rateOUT.png
Attachment 3: 2022-03-02_20-38-16-XYtotal.png
2022-03-02_20-38-16-XYtotal.png
Attachment 4: 2022-03-02_20-39-02-Extotal.png
2022-03-02_20-39-02-Extotal.png
Attachment 5: 2022-03-02_20-42-28-ADCitems.png
2022-03-02_20-42-28-ADCitems.png
Attachment 6: 2022-03-02_20-42-02-goodEvents.png
2022-03-02_20-42-02-goodEvents.png
Attachment 7: 2022-03-02_20-41-36-correlation8.png
2022-03-02_20-41-36-correlation8.png
Attachment 8: 2022-03-02_20-48-27-noise.png
2022-03-02_20-48-27-noise.png
Attachment 9: 2022-03-02_21-24-40-ExTotalZoom.png
2022-03-02_21-24-40-ExTotalZoom.png
Attachment 10: 2022-03-02_21-24-07-EYTotalZoom.png
2022-03-02_21-24-07-EYTotalZoom.png
  365   Wed Mar 2 17:21:35 2022 JM, PMCurrent status

18:21

Temperature setpoint = 113k

Target density OK S1 varies between 2E-8 and 2.5E-8 mbar, E4 between 1E-9 and 5E-9 mbar. (slightly lower than previous check)

Detectors moving well between 23.1mm and 67.43 mm, Move in and move out counters the same. Correlation scalar is on aida1

Rutherford peak varies between 10^3 and 10^2 depending on detector position. No halo peak in rate. Noise peak varies from ~100 peak height when fully in to ~400 peak height when fully out.

Beam intensity ~ 1E8

Merger OK

Data rate ~ 1800 kb/sec

Leakage current ~3.21 uA

Temperature OK
System checks  OK - wr fails

         Base         Current     Difference
aida01 fault      0xab60 :      0xaba3 :      67  
aida02 fault      0x88d3 :      0x8914 :      65  
aida03 fault      0x23aa :      0x23ed :      67  
aida04 fault      0x15d7 :      0x161a :      67  
aida05 fault      0x2f03 :      0x2f20 :      29  
White Rabbit error counter test result: Passed 0, Failed 5

Understand the status reports as follows:-
Status bit 3 : White Rabbit decoder detected an error in the received data
Status bit 2 : Firmware registered WR error, no reload of Timestamp
Status bit 0 : White Rabbit decoder reports uncertain of Timestamp information from WR

Avg ring pressure = 4.4 E-10 mbar

19:20

Temperature setpoint = 113k

Target density OK S1 varies between 2E-8 and 3E-8 mbar, E4 between 1E-9 and 1E-8 mbar. (slightly higher than previous check)

Detectors moving well between 23.1mm and 67.43 mm, Move in and move out counters the same. Correlation scalar is on aida1

Rutherford peak varies between 10^3 and 10^2 depending on detector position. No halo peak in rate.

Beam intensity ~ 1E8

Merger OK

Data rate ~ 1800 kb/sec

Leakage current ~3.20 uA

Temperature OK
System checks  OK - wr fails

         Base         Current     Difference
aida01 fault      0xab60 :      0xaba4 :      68  
aida02 fault      0x88d3 :      0x8915 :      66  
aida03 fault      0x23aa :      0x23ee :      68  
aida04 fault      0x15d7 :      0x161b :      68  
aida05 fault      0x2f03 :      0x2f20 :      29  
White Rabbit error counter test result: Passed 0, Failed 5

Understand the status reports as follows:-
Status bit 3 : White Rabbit decoder detected an error in the received data
Status bit 2 : Firmware registered WR error, no reload of Timestamp
Status bit 0 : White Rabbit decoder reports uncertain of Timestamp information from WR

Avg ring pressure = 4.3 E-10 mbar

Statistics similar to previous check.

Attachment 1: 2022-03-02_18-28-37-Exrate.png
2022-03-02_18-28-37-Exrate.png
Attachment 2: 2022-03-02_18-29-38-Xytotal.png
2022-03-02_18-29-38-Xytotal.png
Attachment 3: 2022-03-02_18-30-33-ExTotal.png
2022-03-02_18-30-33-ExTotal.png
Attachment 4: 2022-03-02_18-32-51-correlation8.png
2022-03-02_18-32-51-correlation8.png
Attachment 5: 2022-03-02_18-31-50-GoodEvents.png
2022-03-02_18-31-50-GoodEvents.png
Attachment 6: 2022-03-02_18-31-24-ADCitems.png
2022-03-02_18-31-24-ADCitems.png
Attachment 7: 2022-03-02_19-48-56-rateIN.png
2022-03-02_19-48-56-rateIN.png
Attachment 8: 2022-03-02_19-50-51-rateOUT.png
2022-03-02_19-50-51-rateOUT.png
  364   Wed Mar 2 14:15:16 2022 JM, PMCurrent status

15:15

Temperature setpoint = 113 K

Target density is OK for now. S1 vaies between 2E-8 and 3.5E-8 mbar. E4 between 1E-9 and 1E-8 mbar.

Detector motors functioning well. Moving between 23.1 mm and 67.43mm. Move in and move out counters the same. Correlation scalar 8 for detectors moving is seen on aida1

Rutherford rate peak varies between 10^3 peak height when fully in and ~10^2 when fully out. No peak from halo seen in rate.

Beam intensity ~ 1E8

Merger OK

Data rate ~ 1700 kb/sec

Run is Run 87 - started 10:34

Data spy restarted at 15:00 so only include Run87 data.

Leakage current = 3.205 uA

Fee Temperatures OK

System checks OK - wr fails

         Base         Current     Difference
aida01 fault      0xab60 :      0xaba3 :      67  
aida02 fault      0x88d3 :      0x8914 :      65  
aida03 fault      0x23aa :      0x23ed :      67  
aida04 fault      0x15d7 :      0x161a :      67  
aida05 fault      0x2f03 :      0x2f20 :      29  
White Rabbit error counter test result: Passed 0, Failed 5

Understand the status reports as follows:-
Status bit 3 : White Rabbit decoder detected an error in the received data
Status bit 2 : Firmware registered WR error, no reload of Timestamp
Status bit 0 : White Rabbit decoder reports uncertain of Timestamp information from WR

Avg ring Pressure = 5.5E-10 mbar

16:45

Temperature setpoint = 113 K

Target density is OK for now. S1 vaies between 2E-8 and 3E-8 mbar. E4 between 1E-9 and 1E-8 mbar.

Detector motors functioning well. Moving between 23.1 mm and 67.43mm. Move in and move out counters the same. Correlation scalar 8 for detectors moving is seen on aida1

Rutherford rate peak varies between 10^3 peak height when fully in and ~10^2 when fully out. No peak from halo seen in rate.

Beam intensity ~ 1E8

Merger OK

Data rate ~ 1300 kb/sec

Temperatures OK

Leakage current ~3.205 uA

System chacks OK - wr fail

         Base         Current     Difference
aida01 fault      0xab60 :      0xaba3 :      67  
aida02 fault      0x88d3 :      0x8914 :      65  
aida03 fault      0x23aa :      0x23ed :      67  
aida04 fault      0x15d7 :      0x161a :      67  
aida05 fault      0x2f03 :      0x2f20 :      29  
White Rabbit error counter test result: Passed 0, Failed 5

Understand the status reports as follows:-
Status bit 3 : White Rabbit decoder detected an error in the received data
Status bit 2 : Firmware registered WR error, no reload of Timestamp
Status bit 0 : White Rabbit decoder reports uncertain of Timestamp information from WR

Avg ring pressure ~ 4.4 E-10 mbar

Statistics similar to previous check.

 

Attachment 1: 2022-03-02_15-23-02-ExRate.png
2022-03-02_15-23-02-ExRate.png
Attachment 2: 2022-03-02_15-24-14-XYtotal.png
2022-03-02_15-24-14-XYtotal.png
Attachment 3: 2022-03-02_15-28-04-correlation8.png
2022-03-02_15-28-04-correlation8.png
Attachment 4: 2022-03-02_15-27-13-ADCitems.png
2022-03-02_15-27-13-ADCitems.png
Attachment 5: 2022-03-02_15-26-34-GoodEvents.png
2022-03-02_15-26-34-GoodEvents.png
Attachment 6: 2022-03-02_16-51-19-aida1-4-stats.png
2022-03-02_16-51-19-aida1-4-stats.png
  363   Wed Mar 2 09:38:50 2022 Spillmann, Uweion source restarted

 

detector position

out: 23.1mm

in: 67.4mm

 

tape: next file

 

gas jet S1: 2.0-2.2 ** -8

 

  362   Wed Mar 2 06:00:49 2022 Spillmann, Uwewaiting

07:00
Target density is hovering between 2.0e-8 and 2.3e-8

ELOG V3.1.4-unknown